OpenStudy (anonymous):

PLEASE HELP! The speed that a car is going on a dry road when the driver puts on the breaks can be estimated by the equation, s = sqrt{24d} where d is the distance that the car skids in feet. A car that was in an accident skidded 148 ft. Find the speed, in miles per hour, that a car was going before the accident occurred. Round your answer to the nearest tenth.

OpenStudy (anonymous):

substitute 148 for d, and evaluate =:\[\sqrt{24*148}=2*\sqrt{3*2*2*2*37}=4\sqrt{222}\]

OpenStudy (anonymous):

how do i evalute that.?

OpenStudy (anonymous):

a calculator :) all I could do was write it the way I did up there in my answer

OpenStudy (anonymous):

my calculator gives me 59.6 mph

OpenStudy (anonymous):

thats what i got lol :) thank you.